OpenStudy (anonymous):

Of the 49 students in a math class, 30 have graphing calculators. 23 of those with graphing calculators were girls, and 7 of those without graphing calculators were girls. How many boys in the class did not have graphing calculators?

OpenStudy (gabbyalicorn):

first, do you know what the operation is?

OpenStudy (anonymous):

Um no?

OpenStudy (anonymous):

12 boys don't have calculators

OpenStudy (shaik0124):

out of 30 graphing calculators 23 of this calculators were girls so remaining is 30-23 =7 boys have graphing calculators

OpenStudy (shaik0124):

now we are left with 49-30 =19 students in this 19 students 7 were girls without graphing calculators so 19-7=12 which are not graphing calculators(boys)

OpenStudy (shaik0124):

did u understand?